OpenStudy (anonymous):

Choose the correct GCF for the following binomial: 6a^4b^2 + 9a^3^b

OpenStudy (anonymous):

r u sure about the second expression?

OpenStudy (anonymous):

6a^4b^2 + 9a^3b oops lol here ya go

OpenStudy (anonymous):

Alright, best way is to do it in bits. So, GCF of 6 and 9?

OpenStudy (anonymous):

3

OpenStudy (anonymous):

Correct, now a^4 and a^3...

OpenStudy (anonymous):

they dont have a GCF

OpenStudy (anonymous):

Well you can divide each of them by a, right?

OpenStudy (anonymous):

No i dont think so i am stuch on this one

OpenStudy (anonymous):

a^4 just means a*a*a*a and to divide by a then you just knock one off to give a*a*a or a^3 So a^4/a would be a^3 In same way, you can divide by a^2, just knock 2 off So a^4/a^2 = a^2 So what would it be if you divided a^4 by a^3?

OpenStudy (anonymous):

its just a right? :)

OpenStudy (anonymous):

Right.

OpenStudy (anonymous):

So now we can go go back, GCF of a^4 and a^3 ?

OpenStudy (anonymous):

so far it is 3a

OpenStudy (anonymous):

a^3, yes.

OpenStudy (anonymous):

And finally, b^2 and b....?

OpenStudy (anonymous):

just b? or is is B^1

OpenStudy (anonymous):

Correct, b= b^1 so you just leave out the ^1 part. So now put the 3 bits together and you have you GCF.

OpenStudy (anonymous):

Factor completely: 64a2b3 − 40ab and with this one the answer is 4(8b2 − 10ab)

OpenStudy (anonymous):

This is a similar idea, you want to find the biggest bit you can "take out of" both parts to put in front of the bracket. So, start with the numbers, GCF of 64 and 40?

OpenStudy (anonymous):

8?

OpenStudy (anonymous):

Right. a^2 and a? b^3 and b?

OpenStudy (anonymous):

a and b^2

OpenStudy (anonymous):

a is right. But b^2 wont divide into b, will it? So that is just b as well. So you can take out 8ab( ) What goes in the brackets?

OpenStudy (anonymous):

Once you done about 20 of these you will just do it without even thinking...:-)

OpenStudy (anonymous):

r u OK now?

OpenStudy (anonymous):

(8ab^2 − 5)?